OpenStudy (anonymous):

Sarah went to an electronics store, purchased 3 CDs and 5 DVDs, and spent $137. Michelle went to the same store, bought 2 CDs and 7 DVDs, and spent $161. How much does each CD sell for at the store? A. $18 B. $13 C. $19 D. $14

OpenStudy (anonymous):

a=cd b=dvd 3a+5b=137 2a+7b=161 2(3a+5b=137) -3(2a+7b=161) 6a+10b=274 -6a-21b=-483 ------------------------ -11b==-209 b=19 each dvd = $19 6a+190=274 6a=84 a=14 each cd = $ 14

OpenStudy (anonymous):

D

OpenStudy (sriram):

3x+5y=137 2x+7y=161 where x is

OpenStudy (sriram):

god u r so fast!!

OpenStudy (anonymous):

it's D-$14

OpenStudy (anonymous):

14

OpenStudy (anonymous):

let price of CD=x price of DVD=y 3x+5y=137 2x+7y=161 ---------- 2*(3x+5y=137) -3*(2x+7y=161) -------------------- 6x+10y=274 -6x-21y=-483 ------------------------ -11y==-209 y=19 price of DVD = $19 plug value of y in eqn 6x+10y=274 6x+190=274 6x=84 x=14 price of CD = $ 14